OpenStudy (anonymous):

a cubic polynomial function f is defined by f(x)=4x^3 +ax^2 +bx +k where a, b, and k are constants. the function f has a local min at x=-1 and the graph of f has a point of inflection at x=-2. find the values of a and b. if the integral from [0, 1] f(x)dx=32, what is the value of k? please help and show all steps!

OpenStudy (turingtest):

the first derivative is\[f'(x)=12x^2+2ax+b\]a minimum here means that\[f'(-1)=12-2a+b=0\]the second derivative is\[f''(x)=24x+2a\]an inflection point here means that\[f''(-2)=-48+2a=0\]use this expression to find a, then use that value of a to find b from f'(-1) then simply do the integral and solve for k